One number is one more than twice another number. The sum of the two squares of the two numbers is 97. I need to find the numbers... can you help?
Answer by mananth(16949)

LET ONENUMBER BE X
THE OTHER NUMBER = 2X+1
...
x^2+(2x+1)^2=97
x^2+4x^2+4x+1=97
5x^2+4x-96=0
5x^2+4x-96=0
5x^2-20x +24x-96=0
5x(x-4)+24(x-4)=0
(x-4)(5x+24)=0
x=4 OR -24/5
...
4 &9
-4.8, -8.6